At a certain temprature and 760ml pressure the volume of gas is 74 ml. If at the same temperature,pressure changed to 740mm. whatis the change in volume?

Grade:11

1 Answers

asit kumar
6 years ago
we can apply boyle`s law i.e. pv=constant. so, 760*74=740*v. then v(final volume)=76ml. so here  the change in volume is 76-74=2ml.